OpenStudy (anonymous):

What is the slope of the line? A. -1/7 B. 1/7 C. –1 D. 1

OpenStudy (marziman):

Can you post the picture?

OpenStudy (anonymous):

@rebeccaxhawaii aye

OpenStudy (math&ing001):

Use the formula : \[ m = \frac{y_2-y_1}{x_2-x_1} \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

-1/7?

OpenStudy (math&ing001):

Not quite, can you show the steps ?

OpenStudy (kiwibyrd):

rise/run. put the number of units it goes up or down over how many units it goes right

OpenStudy (anonymous):

woops 1/7

OpenStudy (math&ing001):

\[ m = \frac{y_2-y_1}{x_2-x_1} = \frac{4-(-3)}{-5-2} = \frac{7}{-7} = -1 \]